# Generative Trajectory Stitching through Diffusion Composition

### v1

@@ 2 Related Work
Diffusion Models For Planning. Many works have studied the applications of diffusion models [57, 23] for generative planning [25, 1, 53, 70, 21, 63, 41, 84]. Diffusion planning has been widely applied in various fields, such as motion planning [5, 42], procedure planning [66], task planning [74, 15], autonomous driving [73, 36], reasoning [77], and reward learning [48]. Many techniques have also been combined with diffusion planning, including hierarchical planning [34, 7, 44], self-evolving planner [35], preference alignment [9], tree branch-pruning [16], refining [30], replanning [82], uncertainty-aware planning [59], equivariance [4]. However, these works are usually constrained to plan within similar horizons as training data. Our work instead proposes a compositional diffusion planning approach that generalizes to much longer horizon via generative trajectory stitching.
@@
Figure 2: Illustrating the Trajectory Stitching Process. Given an unseen start and goal, CompDiffuser generates a long-horizon plan by progressively denoising three trajectory chunks in parallel, with each chunk conditioning on its neighbors to ensure smooth transitions.
@@ 3.1 Compositional Trajectory Modeling
$p_{\theta}(\tau|q_{s},q_{g})\>\propto\>p_{1}(\tau_{1}|\tau_{2},q_{s})~{}p_{K}(\tau_{K}|\tau_{K-1},q_{g})~{}{\prod_{k=2}^{K-1}}p_{k}(\tau_{k}|\tau_{k-1},\tau_{k+1}).$ (2)
@@ 3.2 Training Compositional Trajectory Models
$\mathcal{L}_{\text{start}}=\mathbb{E}_{\tau\in D,t,k}\left[\|\epsilon-\epsilon_{\theta}(\tau_{0}^{t},t~{}|~{}q_{s},\tau_{2}^{t})\|^{2}\right],$ (4)
@@ 3.3 Compositional Trajectory Planning
Parallel Sampling. Our first sampling approach conditions denoising on noisy adjacent trajectory chunk values from the previous denoising timestep, where the update rule is
@@
where $\alpha^{t}$ and $\beta^{t}$ are diffusion specific hyperparameters. This approach allows us to run, in parallel, denoising on each trajectory chunk, as each denoising update only requires the value of the adjacent trajectory chunks at a previous noise level. However, information propagation between the values of adjacent trajectory chunks is limited at each denoising timestep, as each trajectory chunk is denoised independently of the denoising updates of other trajectories.
Autoregressive Sampling. To better couple the values of adjacent trajectory chunks, we propose to decode and denoise each trajectory chunk autoregressively dependent on the values of neighboring chunks at each denoising timestep. In particular, we iteratively denoising each trajectory $\tau_{k}^{t-1}$ starting from the $\tau_{1}^{t-1}$, and condition the denoising of $\tau_{k}^{t-1}$ on the previously decoded value of chunk $\tau_{k-1}^{t-1}$ and the previous timestep decoded value for the future chunk $\tau_{k+1}^{t}$, giving us the sampling equation
@@
This sequential generation process enables stronger coordination among the chunks since each chunk is conditioned on the less noisy versions of previous chunks. However, it requires generating chunks one at a time rather than simultaneously, making it computationally less efficient than parallel sampling. We compare the two sampling schemes in Figure 10, where we empirically find autoregressive sampling leads to improved performance. Additionally, we provide sampling time comparison in Table 7. We use this autoregressive sampling procedure throughout the experiments in the paper and illustrate pseudocode for sampling in Algorithm 2. Given such final set of generated chunks $\tau_{1:K}$, we then merge the chunks together to construct a final trajectory $\tau_{\text{comp}}$ by applying exponential trajectory blending to areas where subchunks $\tau_{k}$ overlap (See Appendix E.2 for implementation details).
@@
Algorithm 1 Training CompDiffuser 1:Require: train dataset $D$, diffusion denoiser model $\epsilon_{\theta}(\tau^{t},t~{}|~{}\mathrm{st\_cond},\mathrm{end\_cond})$, number of training steps $N$, diffusion timestep $T$ 2:for  $i$ = $1\to N$  do 3:  $\tau^{0}\sim D$, sample clean trajectory from dataset 4:  $\tau^{t}\leftarrow\text{add noise to~{}}\tau^{0},~{}t\sim T$ following DDPM 5:  $\tau_{1:K}^{t}\leftarrow$ divide $\tau^{t}$ to $K$ overlapping chunks 6:  # Objective for noisy chunk condition 7:  $\mathcal{L}_{\text{nbr}}=\|\epsilon-\epsilon_{\theta}(\tau_{k}^{t},t~{}|~{}\tau_{k-1}^{t},\tau_{k+1}^{t})\|^{2},~{}k\sim[1,K-1]$ 8:  # Objective for start / end state condition 9:  $\mathcal{L}_{\text{start}}=\|\epsilon-\epsilon_{\theta}(\tau_{0}^{t},t~{}|~{}q_{s},\tau_{2}^{t})\|^{2},~{}q_{s}=\tau_{0}^{0}[0]$ 10:  $\mathcal{L}_{\text{end}}=\|\epsilon-\epsilon_{\theta}(\tau_{K}^{t},t~{}|~{}\tau_{K-1}^{t},q_{g})\|^{2},~{}q_{g}=\tau_{K}^{0}[-1]$ 11:  $\mathcal{L}_{\text{all}}=\mathcal{L}_{\text{nbr}}+\mathcal{L}_{\text{start}}+\mathcal{L}_{\text{end}}$ 12:  Backprop to update $\epsilon_{\theta}(.)$ using $\mathcal{L}_{\text{all}}$ 13:end for 14:return $\epsilon_{\theta}(.)$ Algorithm 2 Autoregressive Trajectory Sampling 1:Models: trained diffusion denoiser model $\epsilon_{\theta}(\tau,t~{}|~{}\mathrm{st\_cond},\mathrm{g\_cond})$ 2:Input: start state $q_{\text{s}}$, goal state $q_{\text{g}}$, number of composed trajectories $K$ 3:Initialize $K$ trajectories $\tau_{1:K}\sim\mathcal{N}(0,I)$ 4:for  $t$ = $T\to 1$  do 5:  # Denoise $\tau_{1}$ conditioned on $q_{s}$ and $\tau_{2}^{t}$ 6:  $\tau_{1}^{t-1}=\epsilon_{\theta}(\tau_{1}^{t},t~{}|~{}q_{\text{s}},\tau_{2}^{t})$ 7:  # Denoise intermediate trajectories $\tau_{2}~{}\text{to}~{}\tau_{K-1}$ 8:  for  $k$ = $12\to K-$ 1  do 9:    $\tau_{k}^{t-1}=\epsilon_{\theta}(\tau_{k}^{t},t~{}|~{}\tau_{k-1}^{t-1},\tau_{k+1}^{t})$ 10:  end for 11:  # Denoise $\tau_{K}$ conditioned on $\tau_{K-1}^{t-1}$ and $q_{s}$ 12:  $\tau_{K}^{t-1}=\epsilon_{\theta}(\tau_{K}^{t},t~{}|~{}\tau_{K-1}^{t-1},q_{\text{g}})$ 13:end for 14:$\tau_{\text{comp}}$ = Merge the denoised trajectories $\tau_{1:K}^{0}$ 15:return $\tau_{\text{comp}}$
@@ 4.1 PointMaze
Figure 4: Qualitative Comparison of DD, GSC and CompDiffuser on OGBench PointMaze Giant. Effective bidirection information propagation enables CompDiffuser to successfully synthesize trajectories from start (bottom left) to goal (upper right), while other methods generate o.o.d trajectories that are disconnected with the start/goal or passing through the walls.
@@ 4.2 High Dimension Tasks
Figure 6: Qualitative Plan generated by CompDiffuser in OGBench AntSoccer Medium. The initial position of the ant is shown by the blue circle and the goal is to move the ball to the pink circle. We plot each individual trajectory $\tau_{1:6}$ separately (as shown from left to right) and mark the ball with a yellow circle for clearer view. These generated trajectories will be merged to form a long-horizon plan $\tau_{\text{comp}}$. Note that our model is only trained on two different types of short trajectories: 1) the ant moves without dribbling the ball; 2) the ant moves while dribbling the ball. At test time, the proposed compositional sampling method can stitch these two types of trajectories and construct end-to-end plans of much longer horizon to solve more difficult tasks – first move to the ball from the far side and then dribble the ball to the goal position.
@@ 4.3 Ablation Studies
Parallel vs. Autoregressive Sampling. We compare the performance of the two proposed compositional sampling schemes, parallel and autoregressive, as presented in Figure 10. Autoregressive sampling consistently outperforms the parallel sampling across various tasks in plan quality, showing that the causal information flow, where each trajectory chunk is conditioned on the already-denoised versions of previous chunks, leads to more coherent and physically consistent transitions between chunks. We include additional discussion and sampling time comparison in Appendix B.3.
@@ 5 Discussion and Conclusion
Limitations. Stitching short trajectories to solve for unseen longer-horizon plans is a challenging problem. While our method excels in our empirical evaluation, there still remain a number of future venues of research. When composing a large sequence of trajectories, the error accumulation in the long chain of bidirectional information propagation might lead to infeasible plans. This issue can be potentially mitigated by using domain/task specific rejection sampling techniques to select from multiple candidate plans or leveraging more expensive MCMC sampling. Moreover, the optimal number of chunks $N$ is task-dependent. Our future research will explore ways to automatically identify a suitable number of chunks $K$, potentially by incrementally increasing the number of chunks dependent on the quality of the generated plan.
@@ E.1 Our Conditional Diffusion Model
Table 10: Number of Trajectories Composed for Each Evaluation Environment. Our diffusion models are trained with a short horizon as listed in the Single Model Horizon column and in test time, we compositionally generate multiple such short trajectories to enable trajectory stitching and construct plans of much longer horizon.

### v2

@@ 2 Related Work
Diffusion Models For Planning. Many works have studied the applications of diffusion models [58, 24] for generative planning [26, 1, 54, 72, 22, 64, 42, 86, 7]. Diffusion planning has been widely applied in various fields, such as motion planning [5, 43], procedure planning [67], task planning [76, 16], autonomous driving [75, 37, 68], reasoning [79], and reward learning [49]. Many techniques have also been combined with diffusion planning, including hierarchical planning [35, 8, 45], self-evolving planner [36], preference alignment [10], tree branch-pruning [17], refining [31], replanning [84], uncertainty-aware planning [60], equivariance [4]. However, these works are usually constrained to plan within similar horizons as training data. Our work instead proposes a compositional diffusion planning approach that generalizes to much longer horizon via generative trajectory stitching.
@@
Figure 2: Illustrating the Trajectory Stitching Process. Given an unseen start (blue circle) and goal (green star), CompDiffuser generates a long-horizon plan by progressively denoising three trajectory chunks in parallel, with each chunk conditioning on its neighbors to ensure smooth transitions.
@@ 3.1 Compositional Trajectory Modeling
$p_{\theta}(\tau|q_{s},q_{g})\>\propto\>p_{1}(\tau_{1}|q_{s},\tau_{2})~{}p_{K}(\tau_{K}|\tau_{K-1},q_{g})~{}{\prod_{k=2}^{K-1}}p_{k}(\tau_{k}|\tau_{k-1},\tau_{k+1}).$ (2)
@@ 3.2 Training Compositional Trajectory Models
$\mathcal{L}_{\text{start}}=\mathbb{E}_{\tau\in D,t,k}\left[\|\epsilon-\epsilon_{\theta}(\tau_{1}^{t},t~{}|~{}q_{s},\tau_{2}^{t})\|^{2}\right],$ (4)
@@ 3.3 Compositional Trajectory Planning
Parallel Sampling. Our first sampling approach conditions denoising on the values of the noisy adjacent trajectory chunks from the previous denoising timestep, where the update rule is
@@
where $\alpha^{t}$ and $\beta^{t}$ are diffusion specific hyperparameters. This approach allows us to run denoising on each trajectory chunk in parallel, as each denoising update only requires the values of the adjacent trajectory chunks at a previous noise level. However, information propagation between the values of adjacent trajectory chunks is limited at each denoising timestep, as each trajectory chunk is denoised independently of the denoising updates of other trajectory chunks.
Autoregressive Sampling. To better couple the values of adjacent trajectory chunks, we propose to denoise each trajectory chunk autoregressively dependent on the values of neighboring chunks at each denoising timestep. In particular, we iteratively denoise each trajectory $\tau_{1:K}^{t}$ starting from the $\tau_{1}^{t}$, and condition the denoising of $\tau_{k}^{t}$ on the previously decoded chunk $\tau_{k-1}^{t-1}$ at the current noise level $t-1$ and the future chunk $\tau_{k+1}^{t}$ at the previous noise level $t$, giving us the sampling equation
@@
This sequential generation process enables stronger coordination among the chunks since each chunk is conditioned on the less noisy version of its previous chunk. However, it requires generating chunks one at a time rather than simultaneously, making it computationally less efficient than parallel sampling. We compare the two sampling schemes in Figure 10, where we empirically find autoregressive sampling leads to improved performance. Additionally, we provide sampling time comparison in Table 7. We use this autoregressive sampling procedure throughout the experiments in the paper and illustrate pseudocode for sampling in Algorithm 2. Given such final set of generated chunks $\tau_{1:K}$, we then merge the chunks together to construct a final trajectory $\tau_{\text{comp}}$ by applying exponential trajectory blending to areas where subchunks $\tau_{k}$ overlap (See Appendix E.2 for implementation details).
@@
Algorithm 1 Training CompDiffuser 1:Require: train dataset $D$, diffusion denoiser model $\epsilon_{\theta}(\tau^{t},t~{}|~{}\mathrm{st\_cond},\mathrm{end\_cond})$, number of training steps $N$, diffusion timestep $T$ 2:for  $i$ = $1\to N$  do 3:  $\tau^{0}\sim D$, sample clean trajectory from dataset 4:  $\tau_{1:K}^{0}\leftarrow$ divide $\tau^{0}$ to $K$ overlapping chunks 5:  $\tau_{1:K}^{t}\leftarrow\text{add noise to~{}}\tau^{0}_{1:K}$, $t\sim T$, following DDPM 6:  # Objective for noisy chunk condition 7:  $\mathcal{L}_{\text{nbr}}=\|\epsilon-\epsilon_{\theta}(\tau_{k}^{t},t~{}|~{}\tau_{k-1}^{t},\tau_{k+1}^{t})\|^{2},~{}k\sim[2,K-1]$ 8:  # Objective for start / end state condition 9:  $\mathcal{L}_{\text{start}}=\|\epsilon-\epsilon_{\theta}(\tau_{1}^{t},t~{}|~{}q_{s},\tau_{2}^{t})\|^{2},~{}q_{s}=\tau_{1}^{0}[0]$ 10:  $\mathcal{L}_{\text{end}}=\|\epsilon-\epsilon_{\theta}(\tau_{K}^{t},t~{}|~{}\tau_{K-1}^{t},q_{g})\|^{2},~{}q_{g}=\tau_{K}^{0}[-1]$ 11:  $\mathcal{L}_{\text{all}}=\mathcal{L}_{\text{nbr}}+\mathcal{L}_{\text{start}}+\mathcal{L}_{\text{end}}$ 12:  Backprop to update $\epsilon_{\theta}(.)$ using $\mathcal{L}_{\text{all}}$ 13:end for 14:return $\epsilon_{\theta}(.)$ Algorithm 2 Autoregressive Trajectory Sampling 1:Models: trained diffusion denoiser model $\epsilon_{\theta}(\tau,t~{}|~{}\mathrm{st\_cond},\mathrm{g\_cond})$ 2:Input: start state $q_{\text{s}}$, goal state $q_{\text{g}}$, number of composed trajectories $K$ 3:Initialize $K$ trajectories $\tau_{1:K}\sim\mathcal{N}(0,I)$ 4:for  $t$ = $T\to 1$  do 5:  # Denoise $\tau_{1}$ conditioned on $q_{s}$ and $\tau_{2}^{t}$ 6:  $\tau_{1}^{t-1}=\epsilon_{\theta}(\tau_{1}^{t},t~{}|~{}q_{\text{s}},\tau_{2}^{t})$ 7:  # Denoise intermediate trajectories $\tau_{2}~{}\text{to}~{}\tau_{K-1}$ 8:  for  $k$ = $12\to K-$ 1  do 9:    $\tau_{k}^{t-1}=\epsilon_{\theta}(\tau_{k}^{t},t~{}|~{}\tau_{k-1}^{t-1},\tau_{k+1}^{t})$ 10:  end for 11:  # Denoise $\tau_{K}$ conditioned on $\tau_{K-1}^{t-1}$ and $q_{g}$ 12:  $\tau_{K}^{t-1}=\epsilon_{\theta}(\tau_{K}^{t},t~{}|~{}\tau_{K-1}^{t-1},q_{\text{g}})$ 13:end for 14:$\tau_{\text{comp}}$ = Merge the denoised trajectories $\tau_{1:K}^{0}$ 15:return $\tau_{\text{comp}}$
@@ 4.1 PointMaze
Figure 4: Qualitative Comparison of DD, GSC and CompDiffuser on OGBench PointMaze Giant. Effective bidirection information propagation enables CompDiffuser to successfully synthesize trajectories from start (bottom left) to goal (upper right), while other methods generate o.o.d trajectories that are disconnected with the start/goal or passing through walls. See Figure 13 for per-segment visualization.
@@ 4.2 High Dimension Tasks
Figure 6: Qualitative Plan generated by CompDiffuser in OGBench AntSoccer Medium Stitch. The initial position of the ant is shown by the blue circle and the goal is to move the ball to the pink circle. We plot each individual trajectory $\tau_{1:6}$ separately (as shown from left to right) and mark the ball with a yellow circle for clearer view. These generated trajectories will be merged to form a long-horizon plan $\tau_{\text{comp}}$. Note that our model is only trained on two different types of short trajectories: 1) the ant moves without dribbling the ball; 2) the ant moves while dribbling the ball. At test time, the proposed compositional sampling method can stitch these two types of trajectories and construct end-to-end plans of longer horizon to solve the more difficult task – first navigate to the ball from the far side and then dribble the ball to the goal position.
@@ 4.3 Ablation Studies
Parallel vs. Autoregressive Sampling. We compare the performance of the two proposed compositional sampling schemes, parallel and autoregressive, as presented in Figure 10. Autoregressive sampling consistently outperforms the parallel sampling across various tasks in plan quality, showing that the causal information flow, where each trajectory chunk is conditioned on the already-denoised (less noisy) version of previous chunk, leads to more coherent and physically consistent transitions between chunks. We include additional discussion and sampling time comparison in Appendix B.3.
@@ 5 Discussion and Conclusion
Limitations. Stitching short trajectories to solve for unseen longer-horizon plans is a challenging problem. While our method excels in our empirical evaluation, there still remain a number of future venues of research. When composing a large sequence of trajectories, the error accumulation in the long chain of bidirectional information propagation might lead to infeasible plans. This issue can be potentially mitigated by using domain/task specific rejection sampling techniques to select from multiple candidate plans or leveraging more expensive MCMC sampling. Moreover, the optimal number of the test-time composed chunks $K$ is task-dependent. Our future research will explore ways to automatically identify a suitable number of chunks $K$, potentially by incrementally increasing the number of chunks dependent on the quality of the generated plan.
@@
Acknowledgments and Disclosure of Funding
We would like to thank Zilai Zeng for helpful early discussion on trajectory stitching and feedback of the manuscript.
@@ E.1 Our Conditional Diffusion Model
Table 10: Number of Composed Trajectories for Each Evaluation Environment. Our diffusion models are trained with a short horizon as listed in the Single Model Horizon column. In test time, we compositionally generate multiple such short trajectories to enable trajectory stitching and construct plans of much longer horizon.